Write the sequence of the following explicit formula:

1. sn = (-4)n , 1 <_ n

2. tn = 92 -5n, 1<_n <_5


1
Expert's answer
2021-02-24T12:19:37-0500

1) Here "S_n=(-4)n" , "n\geq1"

Therefore "S_1=(-4)"

"S_2=(-8)"

"S_3=(-12)"

.

.

"S_n=(-4)n"

"\therefore" In explicit form

{ "(-4),(-8),(-12),.....,(-4)n" }


2)Here "t_n=92-5n,1\leq n\leq5"

Therefore, "t_1=92-5.1=87"

"t_2=92-5.2=82"

"t_3=92-5.3=77"

"t_4=92-5.4=72"

"t_5=92-5.5=67"

"\therefore" In explicit form {"87,82,77,72,67" }